2x - 1 < x*x - 4 < 12
1
Expert's answer
2020-12-24T16:23:57-0500
"\begin{alignedat}{2}
 x^2-2x-3>0 \\
 x^2-16 <0
\end{alignedat}"

"\begin{alignedat}{2}
 (x+1)(x-3)>0 \\
 (x+4)(x-4) <0
\end{alignedat}"

"\begin{alignedat}{2}
 x<-1\ or\ x>3 \\
 -4<x<4
\end{alignedat}"

"x\in(-4,-1)\cup(3,4)"